### v2

# Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization

Abstract
The rapid advancement of Large Language Models (LLMs) highlights the urgent need for evolving evaluation methodologies that keep pace with improvements in language comprehension and information processing. However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios. Our study examines temporal generalization, which includes the ability to understand, predict, and generate text relevant to past, present, and future contexts, revealing significant temporal biases in LLMs. We propose an evaluation framework, for dynamically generating benchmarks from recent real-world predictions. Experiments demonstrate that LLMs struggle with temporal generalization, showing performance decline over time. These findings highlight the necessity for improved training and updating processes to enhance adaptability and reduce biases. Our code, dataset and benchmark are available at https://github.com/FreedomIntelligence/FreshBench.
1 Introduction
Proprietary large language models (LLMs) such as GPT-4 [1], Bard [2], Claude [3], and PaLM [4] has been paralleled by efforts in the open-source community to democratize LLMs. Meanwhile, the rapid push for LLM democratization [5] emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities. Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges. Knowledge-based assessments, exemplified by MMLU [6] and C-Eval [7], are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage. Open-dialogue evaluations like MT-Bench [8] and Alpaca-Eval [9], rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively. Recently, several methods have been proposed to evaluate models at different times and in various contexts[10, 11, 12]. TimeQA[13] evaluates a model’s ability to handle temporal evolution, while Situated QA[14] focuses on a model’s performance in specific contexts. StreamingQA[15] introduces the concept of continuous information flow, while RealTimeQA[11] and FreshQA[12] emphasize real-time and up-to-date information processing capabilities. However, existing methods may be prone to data leakage during their evaluation processes and typically focus on current or past data, lacking exploration of future event prediction capabilities.
We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases [16]. Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations. This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect. By doing so, the inaccessibility of future content would largely reduce the data contamination issue. Moreover, a retrospective perspective on time could be eneficial for temporal bias analysis for LLMs.
To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts. The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia. We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of "intelligence as compression" in a manner free from data leakage. The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios. In summary, the evaluation at temporal generalization offers two main benefits: it eliminates the possibility of data leakage through temporal boundaries, and its results can be objectively assessed without reliance on subjective interpretations.
Our analysis indicates that using validation loss of new text to evaluate models is mostly effective. However, we observe that while performance on language likelihood is a strong indicator of a model’s linguistic capabilities, it does not always align with specific abilities measured by existing benchmarks, which may represent discrete knowledge or other capacities that are important but not directly related to language modeling.
Our contributions are as follows: 1) We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time. 2) We proposed an benchmark Freshbench to test the capabilities of LLMs in predicting future events. This benchmark is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments. It experimentally assesses performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
2 Conceptualization of Temporal Generalization
Figure 1: Schematic figure of temporal generalization.
In this section, we will conceptualize Temporal Generalization in Sec. 2.1 and further measure it in 2.2. From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec. 2.3. In general, prior to the cutoff time, a LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information). However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
2.1 Temporal Generalization
LLMs face the ongoing challenge of remaining applicability as new text evolves. Consequently, it is crucial for these models to not only understand and generate text based on seen (e.g., past) data but also to anticipate and adapt to future data. This inspires temporal generalization as below.
Definition 2.1.
Temporal Generalization refers to the LLMs’ ability to accurately understand, predict, and generate text relevant to past, present, and future contexts.
This requires the models to integrate and apply knowledge from historical and current data to adapt to the evolution of language and emerging trends. Temporal generalization highlights the challenge of designing models that remain effective and accurate over time, without the need for frequent re-training as new data is introduced. LLMs must maintain a delicate balance between leveraging historical data to understand long-standing language patterns and adapting to new expressions, terminology, and topics as they emerge.
2.2 Assessing Temporal Generalization
Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 or II): prognostication prediction in Sec. 2.2.2.
2.2.1 Case I: Compression Intelligence over Time
In the assessment of compression intelligence, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance. We refine our focus to BPC for a more normalized comparison. Unlike traditional BPC which is specific to character-level language models, our adapted Bits Per Character (BPC) metric offers a comparison between models by normalizing the log likelihood based on character count, irrespective of the tokenization method, allowing for a more equitable assessment of model performance across different languages and tokenization schemes, the BPC is calculated as follows:
$\small BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
[…]
2.2.2 Case II: Prognostication Prediction
LLMs could performs prognostication prediction, as seen in [18] which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform. Further, [19] enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus. This approach enhances model training by incorporating varied and relevant data contexts. Moreover, [20] showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve?. As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?". To ensure objectivity in our predictions, answers are typically formatted as ranges within a prediction market scenario. For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
However, since collecting actual future data is impossible, we adopt a retrospective study approach. We use data available after the event’s question was posed and before the prediction market’s closure, treating this as "future" data for model evaluation. This ongoing, online evaluation objectively assesses the model’s predictive capability. We define prediction accuracy, denoted by $Acc$, as follows:
$\small Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
where $N_{\text{correct}}$ is the number of correct predictions among $N_{\text{total}}$ future prediction questions. See. Sec. 4.1 for the collected future prediction question set over time. This metric evaluates the effectiveness of LLMs in predicting future events. This approach also allows continuous evaluation with the latest data, ensuring the assessment reflects the model’s current capabilities and the evolving information landscape.
2.3 Temporal Bias in a Retrospective View
Temporal biases are multifaceted in previous literature. For instance, temporal bias can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality [21] or user preferences over time in point-of-interest recommendations [22]. Recently, biases have raised more challenges in pre-trained language models [23], particularly focusing on under-represented societies. [24] examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance. In the context of LLMs, temporal bias refers to the preferences of LLMs over time.
There exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future. A visual representation of these biases is provided in Fig. 1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
Definition 2.2.
Nostalgia Bias refers to the over-reliance on historical data when generating text. This bias causes models to reflect perspectives, styles, or information prevalent in past training data, leading to outputs that may not account for recent developments. Essentially, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
Definition 2.3.
Neophilia Bias implies a model’s overemphasis on novelty and future trends, potentially at the expense of established information. This bias could result in predictions that are overly speculative and not well-grounded in historical data or current realities. Though less common due to LLM training on historical datasets, Neophilia Bias could be perceived from prompt structuring or interpretation of outputs that prioritize futuristic concepts over historical accuracy.
Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges. Nostalgia Bias can lead to outdated predictions, while overcorrection towards Neophilia Bias might result in speculative predictions disconnected from past trajectories. Balancing these biases is crucial to leverage historical and contemporary data effectively, ensuring future predictions are informed by past patterns and open to new possibilities. This requires careful adjustment of the model’s training data and integration of up-to-date information.
3 Temporal Generalization through the Lens of Compression Intelligence
Temporal generalization in language models evaluates their ability to maintain performance over time, even as data evolves. This section discusses how various models handle data from different periods, showcasing their performance stability or volatility.
3.1 Temporal Bias
3.1.1 Protocol for Quantitative Measurement
Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers. Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges. We also did some basic data pre-processing, due to space limitation, details in Appendix B. An overview of the data volume and the average entry length from a single crawl is provided in Appendix C.
Data sources of Fresh Corpora We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems. They are not only benchmarks of their respective sizes, ranging from 3 billion to 13 billion parameters, but also exemplify the rapid progression in NLP technology. The models originate from diverse sources, showcasing the collaborative and competitive spirit driving innovation in the industry.
Experimental Settings To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss. The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension. Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
3.1.2 Observation of Temporal Bias in Compression Intelligence
To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time. In text data before release, a negative slope indicates the presence of Neophilia Bias. Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood. Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy. What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  3). For models with lower initial BPC, the decrease trend is less pronounced. Taking arXiv text as an example, BPC decreases over time, LLMs exhibit a Neophilia bias rather than a Nostalgia bias, capturing the present more effectively than the past.
Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Fig. 3, provide insight into how well these models process and understand news content over a span of several years. We note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a Nostalgia bias. This trend may stem from the dynamic and evolving nature of news narratives. Detailed result for various sources are in E
Figure 2: BPC on arXiv text from 2015.9 to 2024.1 (smoothed)
Figure 3: BPC on BBC news text from 2020 to 2023, with three-month intervals (smoothed)
3.2 Temporal Generalization
3.2.1 Protocol for Quantitative Measurement
Definition 3.1.
Temporal Bias Index (TBI) For each model in a certain period, we record $BPC_{i}$ at $n$ time points $t_{i}$, forming $n$ data points $(t_{i},p_{i})$, where $i=1,\ldots,n$. We then fit a linear model:
$\small BPC_{i}=a*t_{i}+b$ (3)
where Temporal Bias Index (TBI) is defined as the coefficient $a$. TBI is crucial for understanding the direction and rate of change in model performance over time.
To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets and evaluated their performance using this index. Specifically, We established a base BPC using data from the six months prior to each model’s release. After that, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC; and assessed the overall trend in model performance using TBI. The specific results are detailed in Appendix D.
Table 1: Correlation between Model Attributes and Their Performance Decay Across Various Tasks on Wikipedia and BBC Benchmarks
Corre Model Release MMLU Human- Social- STEM MMLU GSM8 Ar20c Hella- Trust- Wino- Long-
lation Size Date ities Sciences Other wasg fulQA grande bench
Wiki 0.504 0.267 0.599 0.500 0.594 0.393 0.528 0.473 0.380 0.581 0.200 0.632 0.404
BBC -0.044 0.313 0.384 0.317 0.311 0.186 0.153 0.188 -0.071 0.248 0.485 0.204 0.549
3.2.2 Observation
Models perform stably on Wikipedia but show more variation on BBC. In the Wikipedia dataset, models exhibit stable performance, demonstrating their ability to handle well-structured and formal text. However, larger models in the Wikipedia dataset deteriorate faster (correlation coefficient of 0.504), likely due to overfitting. In contrast, the BBC dataset shows significant variations in model performance, indicating insufficient adaptability to dynamic and informal text. From 2020 to 2024, the overall BPC for BBC texts shows an increasing trend, indicating gradual performance deterioration. Individual models show considerable differences in performance on new data; for instance, the Yi, LLaMA2, and Baichuan2 series adapt relatively well to new texts, while Phi-1.5 and OPT series exhibit poorer generalization (see Appendix E). These findings indicate that in dynamic environments, different models have significantly varying adaptation capabilities.
Positive correlation between model attributes and performance decay. (see Tab. 1). There is a positive correlation between model attributes (such as model size and release date) and performance decay. A positive BPC slope indicates an increase in BPC over time, reflecting performance deterioration. Larger model sizes may lead to rapid performance decay, suggesting a trade-off between capability and flexibility. In dynamic environments (such as the BBC dataset), the impact of model size on performance decay is smaller. Additionally, newer models show higher correlations, meaning that despite their advanced capabilities, they may deteriorate faster. These results emphasize the importance of designing models that can maintain robustness over time and across various data environments. As machine learning applications grow in complexity and scale, temporal generalization will become increasingly critical.
4 Temporal Generalization through the Lens of Prognostication Prediction
Advantages of Prognostication Prediction over Compression Intelligence Compression Intelligence is limited by their inapplicability to closed-source models, where key indicators like logits or losses are inaccessible. This exclusion narrows comparative analysis and hinders a comprehensive understanding of the language technology landscape. Additionally, language likelihood can be biased by models finely tuned to specific formats or domains, skewing overall evaluation and presenting a distorted view of a model’s general performance.
This motivated us to research temporal generalization and bias from the perspective of future prediction. In this section, we investigate tempera generalization (in Sec. 4.1) and biases (Sec. 4.2) through the Lens of Prognostication Prediction.
4.1 Temporal Biases
4.1.1 Protocol for Quantitative Measurement: a Hypothesis Testing Method
To distinguish the past and the future, we take the border as the release date of the model. For example, if a LLM is released at Jan. 2024 (also denoted as ‘cutoff time’ in this work), the time after Jan. 2024 would be the future. We could collect question prediction tasks in different periods. Based on different periods, We compute the prediction accuracy in these different periods, denoted as $acc_{\text{past}}$ and $acc_{\text{future}}$, where accuracy of the present (somewhere between the past and future) the is denoted as $acc_{\text{present}}$.
To investigate the two mentioned temporal biases in Sec. 2.3, we propose two hypothesis testings as follows. These tests help practitioners make informed decisions despite the randomness inherent in the question datasets.
Hypothesis Testing 1.
Neophilia Bias:
$\small T_{1}:\ H_{0}:\mathbf{E}[acc_{\text{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{past}}]<\mathbf{E}[acc_{\text{present}}],\$ (4)
This hypothesis tests if there is a significant increase in model accuracy for current questions compared to past questions, suggesting a bias towards newer content. The p-value $\widehat{p}$ quantifies the strength of evidence against the null hypothesis.
Hypothesis Testing 2.
Nostalgia Bias:
$\small T_{2}:\ H_{0}^{\prime}:\mathbf{E}[acc_{\textrm{past}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}^{\prime}:\mathbf{E}[acc_{\text{past}}]>\mathbf{E}[acc_{\text{present}}],$ (5)
Conversely, this hypothesis tests if there is a significant increase in model accuracy for past questions, indicating a preference for older or historical content. The p-value $\widehat{p}^{\prime}$ quantifies the strength of evidence against the null hypothesis.
P-Value Computation
The p-values for assessing the shifts in accuracy are computed using the standard normal cumulative distribution function $\Phi$ as follows:
$\small\widehat{p}=1-\Phi\left(\frac{acc_{\text{1}}-acc_{\text{2}}}{\sqrt{\frac{acc_{\text{1}}(1-acc_{\text{1}})}{n_{1}}+\frac{acc_{\text{2}}(1-acc_{\text{2}})}{n_{2}}}}\right),$ (6)
$acc_{\text{1}}$ and $acc_{\text{1}}$ could be either $acc_{\textrm{past}}$ or $acc_{\textrm{present}}$, which is based on which Hypothesis Testing is used. Note that the two Hypothesis Testings are symmetrical.
Experimental setting We have sourced all questions from GoodJudgmentOpen [25] through web scraping and API downloads. The collected data includes initiation and conclusion dates, domains, questions, descriptions, and answers, which are ordered appropriately for further processing. Throughout this experiment, approximately 50 GPU hours on NVIDIA A800 were consumed.
We collected 2532 questions from GoodJudgmentOpen  [25], earliest one was proposed on 2015-09-01 and the most recent one was closed on 2024-05-25. It is "Will the adopted EU Artificial Intelligence (AI) Act be published in the Official Journal of the European Union before 25 May 2024?" Most common tags are in Appendix F.
Table 2: Comparison of Model Performance in Different Pre-Release Periods (Compared to 0-20 months before Release), uses asterisks to denote significance levels: (p < 0.05), ††(p < 0.01), and ††† (p < 0.001), $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$ refers to Neophilia Bias, $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$ refers to Nostalgia Bias.
Models 20-40 Months Before Release 40-60 Months Before Release 60-80 Months Before Release
OPT-13B -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
OPT-2.7B -- $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$  --
LLaMA-7B -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
Pythia-12B $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Falcon-rw-1B -- -- --
Baichuan-7B-Chat -- -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$${}^{\dagger}$
LLaMA-2-13B -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
LLaMA-2-7B $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  --
LLaMA-2-7B-Chat -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Zhongjing-Base -- -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
InternLM-Chat-7B $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- --
Baichuan2-7B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- --
Mistral-7B-v0.1 $\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\boldsymbol{\rightarrow}}}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Phi-1.5 $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Baichuan2-13B-Base $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Baichuan2-13B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Colossal-LLaMA-2-7B-Base $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$
Qwen-14B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Qwen-7B -- -- --
Skywork-13B-Base -- -- --
Zephyr-7B-beta $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Yi-6B -- -- --
Yi-6B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$
Qwen-1.8B -- -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
RWKV-v5-Eagle-7B -- -- --
Phi-2 $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$
Command R+ $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Mixtral-8x22B-Instruct-v0.1 -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
Phi-3-mini-4k-instruct -- $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†
Qwen1.5-110B-Chat $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$  $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$†† $\boldsymbol{\boldsymbol{\boldsymbol{\leftarrow}}}$††
4.1.2 Observation
Neophilia Bias is prevalent closer to release, whereas Nostalgia Bias occurs earlier. As shown in Tab. 2, among the biases, Nostalgia is the most common, followed by periods where neither bias is significant (shown as -), with Neophilia Bias being the least frequent.
Phi-1.5 and Baichuan2-13B series show Nostalgia Bias in all three periods before release. Pythia-12B and Mistral-7B-v0.1 show Neophilia Bias in 20 to 40 months before release but show more significant Nostalgia in another two periods, which we attribute to random fluctuations rather than a consistent trend. OPT-2.7B is Neophilia overall, but this observation is likely due to its consistently near-random performance across various periods, rather than a true preference for novelty.
Models tend to be stronger in earlier periods before release, with performance gradually declining over time. Using the -20 to 0 months period as an anchor, if we assign the three periods to represent the near past (20 to 40 months before release), the middle past (40 to 60 months before release), and the distant past (60 to 80 months before release), we observe that the further back in time we go, the more pronounced the Nostalgia Bias becomes, and it is also more significant. This suggests that the models gradually degrade during the pre-release period, rather than experiencing a rapid decline around the cutoff period.
This observation reveals that models generally favor older knowledge and prefer information from more distant past periods. The distribution of knowledge is not uniform, nor does it tend to be closer to the present. This poses a potential risk in our utilization of these models because there is an expectation that the knowledge utilized should be as up-to-date as possible, ideally close to the model’s release date.
4.2 Temporal Generalization
4.2.1 Protocol
In the prospective view when especially using LLMs for prognostication prediction, there might be a temporal degeneration issue defined as below:
Definition 4.1.
Temporal Degeneration The capability of prognostication prediction might be degraded over time due to the larger generalization gap for longer time.
To qualitatively measure Temporal Degeneration, we propose a new hypothesis testing.
Hypothesis Testing 3.
Temporal Degeneration
$\small T_{3}:$ $\ H_{0}:\mathbf{E}[acc_{\text{future}}]=\mathbf{E}[acc_{\text{present}}]\ \text{versus}\ H_{1}:\mathbf{E}[acc_{\text{future}}]<\mathbf{E}[acc_{\text{present}}],$ (7)
Similar to Sec 4.1.1, $acc_{\text{future}}$ represents the model accuracy on questions proposed after the model’s release. This hypothesis assesses whether there is a significant decline in model accuracy, indicating a degeneration over time. The hypothetical p-value is similar to that of Eq. 6.
Table 3: Accuracy performance of various models across different periods. The presence of asterisks (*) denotes statistically significant decreases in performance during specific periods (p < 0.05 for *, p < 0.01 for ††, and p < 0.001 for ††*), benchmarked against their accuracy before release. Grey cells indicate the models have not been released in this period and represent the average scores of the models prior to their release.
Model 23/1 -23/2 23/3 -23/4 23/5 -23/6 23/7 -23/8 23/9 -23/10 23/11 -23/12 24/1 -24/2 24/3 -24/4
OPT-13B 0.55 0.48 0.31** 0.39 0.41 0.52 0.37* 0.27**
OPT-2.7B 0.17 0.20 0.17* 0.22 0.25 0.22 0.26 0.26
LLaMA-7B 0.21* 0.17** 0.16** 0.22* 0.25 0.22** 0.22* 0.24*
Pythia-12B 0.48 0.47 0.32** 0.43 0.48 0.39* 0.43 0.31**
Falcon-rw-1B 0.22 0.16 0.14* 0.28 0.27 0.21 0.20 0.24
Baichuan-7B-Chat 0.32 0.24* 0.30 0.23 0.19** 0.20* 0.34
InternLM-Chat-7B 0.35 0.23** 0.33 0.27 0.29 0.30 0.31
GPT-3.5-turbo-230613 0.49 0.29** 0.37* 0.45 0.44 0.37* 0.30**
GPT-4-230613 0.6 0.39** 0.43* 0.41** 0.37** 0.48 0.40**
LLaMA-2-13B 0.52 0.43 0.48 0.36** 0.39* 0.40*
LLaMA-2-7B 0.27 0.26 0.27 0.22 0.26 0.34
LLaMA-2-7B-Chat 0.45 0.30* 0.46 0.35** 0.37 0.38
Baichuan2-7B-Chat 0.37 0.35 0.43 0.31 0.37 0.41
Baichuan2-13B-Base 0.49 0.30** 0.46 0.42 0.28** 0.40*
Baichuan2-13B-Chat 0.6 0.33** 0.43 0.41* 0.26** 0.27**
Zhongjing-Base 0.38 0.30 0.33 0.26* 0.39
Mistral-7B-v0.1 0.44 0.41 0.34* 0.28* 0.34*
Phi-1.5 0.41 0.38 0.39 0.30 0.25**
Colossal-LLaMA-2-7B-Base 0.42 0.43 0.37 0.28* 0.33*
Qwen-14B-Chat 0.39 0.25** 0.23** 0.33 0.27**
Qwen-7B 0.36 0.36 0.30 0.33 0.25**
Skywork-13B-Base 0.37 0.39 0.29* 0.30 0.27*
ChatGLM3-6B 0.42 0.38 0.33* 0.26** 0.34
Zephyr-7B-beta 0.4 0.32 0.31* 0.28* 0.28*
Yi-6B 0.29 0.27 0.16** 0.22 0.34
Yi-6B-Chat 0.38 0.25** 0.39 0.35
Qwen-1.8B 0.35 0.22** 0.37 0.30
RWKV-v5-Eagle-7B 0.29 0.29 0.33 0.28
GPT-4-231106 0.66 0.44** 0.39** 0.36**
Phi-2 0.31 0.23* 0.22 0.30
Command R+ 0.54 0.38**
Mixtral-8x22B-Instruct-v0.1 0.56 0.40**
Phi-3-mini-4k-instruct 0.38 0.26**
Qwen1.5-110B-Chat 0.56 0.40**
4.2.2 Observation
Earlier models perform worse. As the latest scores in Table 3 shows, models released earlier perform worse. The poor performance may be due to two factors. One is the natural decay over time, as evidenced by the clear downturn in LLaMA-7B’s performance. Over time, earlier models fail to adapt to new data. The other is the initial lack of advanced training techniques, which may have resulted in models like OPT showing near-random performance from the outset, without significant changes. This is mainly because earlier models were trained with less advanced techniques, leading to poorer performance.
Most models exhibit weak temporal generalization ability. Experimental results show that many models have significant fluctuations in accuracy percentages after release, indicating weak temporal generalization ability. Models with strong predictive capabilities fail to maintain their scores, even if only for a short period. For instance, GPT-4-230613 has demonstrated a significant decline in four out of six periods after release, while GPT-4-231106 has seen a significant decrease across all three observed periods. This variability and instability suggest that models have insufficient adaptive capabilities when dealing with new, unseen data. Some models exhibit near-random performance after the cutoff event, highlighting their limited understanding of the world and inability to generalize effectively from past information to future contexts. This shows that current models’ temporal generalization ability is significantly lacking and needs further improvement.
More advanced models perform better in future predictions. . For instance, Qwen1.5-110B-Chat, Command R+(104B), Mixtral-8x22B-Instruct-v0.1, along with GPT-4, have achieved the highest scores during the period from March to April 2024. It is worth noting that certain versions of LLaMA-2 and Baichuan2 have also performed commendably.
This observation may stem from a bias, as models like Qwen1.5-110B-Chat , Command R+(104B), Mixtral-8x22B-Instruct-v0.1 are the most recently released model, and a more robust evaluation awaits further data updates. The advanced training techniques employed by these models likely contribute to their enhanced accuracy and stability. As time advances and an increasing amount of data is accumulated, the models’ assessments are expected to become more precise and valuable.
5 Conclusion
We rigorously define and qualify temporal generalization and bias within LLMs, uncovering significant challenges in predicting and adapting to future contexts. Our experiments reveal that LLMs often struggle with temporal biases, affecting reliability in dynamic scenarios requiring accurate forecasting. The evaluation framework and datasets we introduce allow for an in-depth assessment of model performance over time. The findings highlight shortcomings in existing benchmarks related to temporal shifts and underscore the urgent need for improvements in LLM training and updating processes to enhance adaptability and mitigate biases.
Limitations
While we strive to collect textual data from a variety of sources such as news articles, scientific papers, and forum discussions, the proportion and diversity of the texts may still fall short of fully reflecting the diversity and complexity of language. Moreover, such evaluations on text loss require access to the model’s logits, which is challenging for closed-source models or models accessed via APIs.
Future Work
Moving forward, it is crucial to refine LLM development and evaluation techniques to better handle evolving informational environments, ensuring their continued relevance and utility in various applications.
Looking ahead, our agenda for future work includes several ambitious goals aimed at enhancing the versatility and depth of our evaluation framework. Key among these objectives are:
  • Integrating a wider array of models, particularly focusing on those based on the Transformer architecture, such as Griffin and Mamba models.
  • Expanding the number of checkpoints accessed for a more comprehensive temporal analysis of model performance.
  • Delving into the relationship between token-level loss metrics and overall scores, which could offer more granular insights into model capabilities.
  • Developing methodologies to assess the obsolescence of benchmarks and models, potentially offering a dynamic approach to evaluating the relevance and efficiency of language models over time.
These initiatives represent our commitment to advancing the state of the art in language model evaluation, ensuring that our tools remain at the forefront of technological progress and innovation.
Community Call to Action
We urge the community to take note of the above phenomenon: models tend to have a better comprehension of earlier events. This observation supports the findings in  [26], which suggest that models are most familiar with knowledge up to 2020. This observation is contrary to our initial hypothesis that models may overfit current events, which is significantly more concerning. It suggests that caution is needed, especially for those hoping to use model judgments directly.
Ethics Statement
We fully acknowledge the need to adhere to ethical standards in handling and generating textual data. Hence, we commit to:
  • We utilize various datasets in our experiments, including publicly available sources like news articles and academic papers, which do not contain personally identifiable information or offensive content. However, some datasets created or used during our study contain sensitive information or are bound by specific usage agreements that prevent public sharing. Therefore, these datasets are utilized strictly within our experiments and are not released as part of our open-source materials, adhering to ethical guidelines and agreements.
  • Responsibility for Content Generated by Open-Source Models: We recognize that content generated by our system or through the use of open-source models involved in the evaluation might reflect or amplify biases. While efforts are made to minimize these biases and improve the fairness of the content, we cannot take responsibility for the generated content. We encourage users and developers to critically engage with these models and remain vigilant of their outputs.
  • Addressing Data Leak Issues: Our system aims to tackle data leak issues by dynamically generating benchmarks and ensuring models generalize to new or future texts. We commit to continuously monitoring and updating our evaluation methods to avoid potential data leakage, ensuring the validity and reliability of our assessment results.
  • Transparency: We are committed to maintaining the highest level of transparency in our work, including our evaluation methods, data sources used, and assessment outcomes. We encourage open critique and suggestions for improvement of our methodology, aiming to foster knowledge sharing and technological advancement.
We believe that through this approach, we can offer a valuable resource to the research and development community, while fostering deep reflection and discussion on ethical issues in the artificial intelligence field.
Licence
5.1 Good Judgment Inc.
For personal use of its services and content, a limited, revocable, non-exclusive license is granted. Without explicit written permission from Good Judgment Inc., this license restricts the collection, aggregation, copying, duplication, display, or derivative use of its content.
5.2 arXiv
The arXiv dataset is licensed under the Creative Commons Zero (CC0 1.0) license. This means it is in the public domain, and users can freely use, modify, and distribute the data without any restrictions.
5.3 Wikipedia
Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike (CC BY-SA) license. This allows users to share and adapt the content as long as appropriate credit is given, and any derived works are distributed under the same license.
Other text will not be released.
Acknowledgement
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing.
Appendix A Testing Statistical Hypothesis
Due to the inherent randomness in data, directly comparing outcomes of two experiments is insufficient to justify the superiority of one outcome over the other. To address this issue, practitioners often resort to hypothesis testing to make convincing decisions, like those demonstrated in [27]. A board application of hypothesis testing lies in the A/B testing literature, which assesses the performance of various features of a product and performs post-experiment inferences, as introduced in [28].
Appendix B Details of Data Collection
The data for our system is sourced from a diverse set of online platforms, each selected for its relevance to different data types and domains. We utilize a Python-based crawling framework, which is detailed in our publicly available repository. This framework is designed to be adaptable, allowing for the inclusion of additional sources as they become relevant.
The primary data sources include:
  • Financial News: Crawled from Yahoo, providing up-to-date information on global financial trends.
  • Political News: BBC News is used to gather the latest political developments worldwide.
  • Discussion Forums: Reddit is scraped to capture current discussions across a variety of tofigures.
  • Online Literature: Wattpad offers a rich source of contemporary fiction and non-fiction.
  • Encyclopedia: Wikipedia’s latest changes feed is monitored for updates across all domains.
  • Academic Papers: arXiv is used to access the newest research across multiple scientific fields.
  • Code Repositories: GitHub Trending provides insights into the latest developments in software engineering.
  • Question and Answer Forums: Quora, focusing on popular topics such as Technology, Mathematics, Health, and Movies.
To facilitate efficient and effective data collection, our system’s environment is configured with a specific Python package index URL and utilizes Playwright for web navigation and content extraction. This setup addresses potential challenges, such as page refresh requirements on Yahoo and hostname verification issues encountered with certain websites.
Our system employs a combination of tools and libraries, including Requests, Playwright, and various PDF processing utilities, to collect the latest textual data from the internet. This approach enables us to dynamically update our benchmark datasets with fresh information, ensuring that our evaluation reflects current language use and information trends.
Data Pre-processing For some raw content clean the text by removing these tags to ensure that only the content text is analyzed.
Filtering out short texts: Extremely short texts may not provide enough contextual information for analysis or model training. Setting a minimum text length threshold of 100 chars to filter out such instances.
External Dataset For the analysis of text likelihood and bias, we employed not only our own collection of historical arXiv data but also sampled entries from two additional datasets. We incorporated data from the BBC News and Wikitext datasets, which are publicly available through the Hugging Face datasets repository. Specifically, the datasets used were bbc_news_alltime and wikitext_alltime, accessible via the following URLs:
Sampling from these datasets was carried out every three months, with each instance involving the collection of 50 entries per dataset. This structured sampling approach was designed to ensure a consistent and representative analysis of content over time, enabling an effective assessment of textual bias and likelihood trends.
Appendix C Data Overview of One Crawl
Tab. 4 presented below offers a comprehensive overview of the diverse sources and categories of texts that were analyzed in one data crawl. Classified into distinct groups such as Academic STEM, Academic Non-STEM, Internet QA, and Internet articles, the table delineates the number of texts extracted from each source along with their average length. This includes detailed counts from different academic archives like arXiv, various categories from Internet question and answer platforms like Reddit and Quora, and articles from well-known online platforms including Wiki and BBC.
Table 4: Overview of text categories with their sources counts, and average lengths in one sample crawl.
Category Source Count Length
Academic STEM arXiv Mathematics 50 42850
arXiv Computer Science 50 43336
arXiv Physics 50 42580
arXiv Statistics 50 42619
arXiv Electrical Engineering and Systems Science 50 43027
Academic Non-STEM arXiv Economics 50 42592
arXiv Quantitative Finance 50 43969
arXiv Quantitative Biology 50 42974
Code GitHub 30 21849
Internet QA Reddit 583 643
Quora Health 20 1150
Quora Movies 69 1349
Quora Mathematics 70 1365
Quora Technology 80 1674
Internet article Wiki 220 4091
Wattpad 38 6615
BBC 45 4792
Yahoo 41 5373
Appendix D Detailed results of BPC generalization on Wiki and BBC dataset
Table 5: Percentage Change in BPC Following Release, measured by BBC news data from 2020 to 2024.3. Dash ("-") symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
OPT-13B 0.450 2.273 -2.087 1.265 0.839 0.572
OPT-2.7B 0.475 1.970 -2.228 1.192 1.143 0.519
LLaMA-7B 0.450 1.270 0.177 -7.776 3.266 -0.766
Pythia-12B 0.480 -0.233 -3.888 -2.612 -1.012 -1.936
Falcon-rw-1B 0.510 0.459 -7.814 1.201 -1.474 -1.907
Baichuan-13B-Chat 0.467 -3.645 -1.078 0.413 - -1.436
LLaMA-2-13B 0.416 -6.928 3.502 1.683 - -0.581
LLaMA-2-7B 0.434 -7.314 3.092 1.279 - -0.981
LLaMA-2-7B-Chat 0.505 -8.308 2.695 0.664 - -1.650
Zhongjing-Base 0.465 -8.488 2.662 -1.290 - -2.372
InternLM-Chat-7B 0.515 -2.476 -1.344 0.236 - -1.195
Baichuan2-7B-Base 0.451 -3.059 0.595 -4.649 - -2.371
Baichuan2-7B-Chat 0.517 -3.571 -0.038 -5.315 - -2.975
Mistral-7B-v0.1 0.398 1.110 5.347 -0.661 - 1.932
Phi-1.5 0.602 -1.772 1.131 -1.837 - -0.826
Baichuan2-13B-Base 0.439 -2.943 0.882 -4.390 - -2.150
Baichuan2-13B-Chat 0.495 -3.541 -0.995 0.232 - -1.435
Colossal-LLaMA-2-7B-Base 0.618 2.825 3.715 - - 3.270
Qwen-14B-Chat 0.445 2.682 4.847 - - 3.764
Qwen-7B 0.441 1.892 3.689 - - 2.791
Qwen-7B-Chat 0.466 1.809 3.606 - - 2.708
Skywork-13B-Base 0.429 7.163 3.261 - - 5.212
ChatGLM3-6B 0.832 5.567 1.684 - - 3.626
Zephyr-7B-beta 0.435 7.340 3.928 - - 5.634
Yi-6B 0.427 6.478 3.144 - - 4.811
Yi-6B-Chat 0.446 3.030 -2.798 - - 0.116
Qwen-1.8B 0.520 1.899 -2.082 - - -0.091
Qwen-1.8B-Chat 0.599 1.635 -2.033 - - -0.199
RWKV-v5-Eagle-7B 0.440 4.956 2.278 - - 3.617
TinyLLaMA-1.1B-Chat-v0.6 0.513 2.539 -3.678 - - -0.569
Phi-2 0.487 1.882 - - - 1.882
Table 6: Percentage Change in Language Likelihood Following Release, measured by wiki data from 2020 to 2024.3. Dash ("-") symbols in the table indicate missing data points due to the later release dates of some models, which means there is no available data for those specific time intervals.
Model Base BPC BPC Change at 3 months (%) BPC Change at 6 months (%) BPC Change at 9 months (%) BPC Change at 12 months (%) mean
OPT-13B 0.075 -0.623 0.105 -0.413 -0.818 -0.437
OPT-2.7B 0.082 -0.658 0.024 -0.528 -0.915 -0.519
LLaMA-7B 0.052 0.214 -0.425 0.379 -0.067 0.025
Pythia-12B 0.071 -0.578 -0.933 -0.859 -0.903 -0.818
Falcon-rw-1B 0.083 -1.272 -1.113 -1.131 -1.207 -1.181
Baichuan-13B-Chat 0.058 0.019 0.124 -0.125 - 0.006
LLaMA-2-13B 0.041 1.700 1.446 1.796 - 1.647
LLaMA-2-7B 0.050 0.774 0.349 0.517 - 0.547
LLaMA-2-7B-Chat 0.060 0.132 -0.186 -0.093 - -0.049
Zhongjing-Base 0.050 0.756 0.269 0.569 - 0.531
InternLM-Chat-7B 0.080 -0.482 -0.301 -0.250 - -0.344
Baichuan2-7B-Base 0.059 0.523 0.157 0.185 - 0.289
Baichuan2-7B-Chat 0.068 0.343 -0.035 -0.073 - 0.078
Mistral-7B-v0.1 0.053 0.381 -0.028 0.018 - 0.124
Phi-1.5 0.098 0.066 -0.125 -0.248 - -0.102
Baichuan2-13B-Base 0.053 1.019 0.797 0.900 - 0.905
Baichuan2-13B-Chat 0.062 0.299 0.628 0.512 - 0.480
Colossal-LLaMA-2-7B-Base 0.065 -0.042 -0.199 - - -0.121
Qwen-14B-Chat 0.049 1.301 1.939 - - 1.620
Qwen-7B 0.061 0.428 0.430 - - 0.429
Qwen-7B-Chat 0.065 0.302 0.226 - - 0.264
Skywork-13B-Base 0.050 0.075 0.276 - - 0.175
ChatGLM3-6B 0.105 -0.187 -0.245 - - -0.216
Zephyr-7B-beta 0.059 -0.049 -0.033 - - -0.041
Yi-6B 0.058 -0.071 0.022 - - -0.024
Yi-6B-Chat 0.060 -0.028 0.023 - - -0.003
Qwen-1.8B 0.080 -0.131 -0.251 - - -0.191
Qwen-1.8B-Chat 0.091 -0.331 -0.496 - - -0.414
RWKV-v5-Eagle-7B 0.065 0.149 0.220 - - 0.184
TinyLLaMA-1.1B-Chat-v0.6 0.070 -0.332 -0.473 - - -0.403
Phi-2 0.072 -0.037 - - - -0.037
Appendix E TBI of Models across the arXiv, BBC, and Wiki datasets
Table 7: Comprehensive list of models with their release dates and metrics for TBI (multiplied by 1000) across the arXiv, BBC, and Wiki datasets. Negative TBI values indicate a trend of increasing BPC, which refers to decreasing performance on MMLU.
Model Name Released Date TBI*1000 (arXiv) TBI*1000 (BBC) TBI*1000 (Wiki)
OPT-13B May 2022 -19.4 1689.0 -30.2
OPT-2.7B May 2022 -18.8 1080.7 -46.9
LLaMA-7B Feb 2023 -15.3 713.0 52.0
Pythia-12B Mar 2023 -8.1 802.7 -11.1
Falcon-rw-1B Apr 2023 -26.8 607.4 -47.7
Baichuan-13B-Chat Jun 2023 -11.7 685.7 4.8
Baichuan-7B-Chat Jun 2023 -14.6 726.6 -16.5
LLaMA-2-13B Jul 2023 -11.1 769.6 65.9
LLaMA-2-7B Jul 2023 -12.6 684.6 23.2
Baichuan-13B-Chat Jul 2023 -7.0 717.2 12.1
Zhongjing-Base Jul 2023 -15.0 759.7 56.9
InternLM-Chat-7B Jul 2023 -19.3 574.0 -22.0
Baichuan2-7B-Base Aug 2023 -9.5 687.5 3.7
Baichuan2-7B-Chat Aug 2023 -8.4 781.8 -11.6
Mistral-7B-v0.1 Sep 2023 -15.1 756.8 10.4
Phi-1.5 Sep 2023 -26.0 1212.9 -60.7
Baichuan2-13B-Base Sep 2023 -9.1 747.9 15.8
Baichuan2-13B-Chat Sep 2023 -7.0 717.2 12.1
Colossal-LLaMA-2-7B-Base Sep 2023 -21.2 696.3 -36.2
Qwen-14B-Chat Sep 2023 -7.9 762.2 51.5
Qwen-7B Sep 2023 -11.7 662.4 7.4
Qwen-7B-Chat Sep 2023 -13.2 705.7 -3.6
Skywork-13B-Base Oct 2023 -16.3 549.1 30.8
ChatGLM3-6B Oct 2023 -25.5 412.6 -70.4
Zephyr-7B-beta Oct 2023 -21.1 506.9 3.7
Yi-6B Nov 2023 -9.4 451.7 24.6
Yi-6B-Chat Nov 2023 -9.4 506.5 20.2
Qwen-1.8B Nov 2023 -22.7 374.0 -46.2
Qwen-1.8B-Chat Nov 2023 -25.3 426.8 -50.3
RWKV-v5-Eagle-7B Nov 2023 -12.8 511.3 4.2
TinyLLaMA-1.1B-Chat-v0.6 Dec 2023 -25.0 300.1 -45.3
Phi-2 Dec 2023 -21.2 669.3 -44.7
Appendix F Most common tags of the Gjo questions
Table 8: Frequency of Topics
Topic Count
Business 648
Non-US Politics 453
Security and Conflict 446
Society 425
US Politics 363
Technology 359
Finance 299
Elections and Referenda 288
Economic Indicators 273
Foreign Policy 271
US Policy 247
Health 245
Leader Entry/Exit 219
Economic Policy 160
Sports 148
Entertainment 148
Environment 80
Open 16
Table 9: Most common tags of the Gjo questions
Appendix G Correlation with Existing Benchmarks
Following our previous chapter I, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in I. We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items. Each group represents a significant area of interest.
In our analysis, we focus on how the correlation between these categories and various benchmarks evolves over time. We use correlation coefficients to measure the strength of association between the performance scores of questions within each category and the benchmarks across different times.
Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment. These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics. Detailed figures are in Appendix N.
Despite relatively high initial correlations in 2 question groups, there is a general lack of correlation between model predictions and existing benchmarks (mean correlation coefficient around zero). In addition, significant deviations occur in nearly half of the cases from early stages to subsequent periods. This suggests that the models’ understanding and alignment with benchmarks can vary substantially over time. So temporal generalization provides novel insights and draws attention to an aspect that has been often ignored by previous benchmarks.
Appendix H Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
(a) STEM: base models
(b) non-STEM: base models
(c) Non-STEM: chat models
Figure 4: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores. Each row represents a specific truncation length, while the columns represent different benchmark score categories. The values in the cells indicate the corresponding correlation coefficients. (a) presents the correlations by loss for all 18 base models on STEM Academia texts. (b) presents the correlations by loss for all 18 base models on non-STEM Academia texts. (c) presents the correlations by loss for all 15 chat models on non-STEM Academia texts.
(a) STEM: base models
(b) non-STEM: base models
(c) Non-STEM: chat models
Figure 5: The heatmaps display the correlations by loss for a group of models across different text chunk lengths and different benchmark scores. Each row represents a specific truncation length, while the columns represent different benchmark score categories. The values in the cells indicate the corresponding correlation coefficients. (a) presents the correlations by loss for all 18 base models on STEM Academia texts. (b) presents the correlations by loss for all 18 base models on non-STEM Academia texts. (c) presents the correlations by loss for all 15 chat models on non-STEM Academia
(a) STEM: all chat models
(b) STEM: 7B level chat models
Figure 6: (a) presents the correlations by loss for all 15 chat models on STEM Academia texts. (b) presents the correlations by loss for all 9 7B level chat models on STEM Academia texts.
From Fig. 6 we can conclude that the scale of models being observed does not significantly alter outcomes, suggesting that performance correlations across benchmarks are generally consistent regardless of model size.
Appendix I Correlation with Existing Benchmarks
The Selected Existing Benchmarks
We explored the correlation between BPC and existing benchmarks, particularly focusing on how specific content types, such as GSM, influence model performance. The selected benchmarks encompass common sense reasoning, depth and breadth of knowledge across diverse subjects, truthfulness, natural language understanding, and mathematical reasoning, ensuring a well-rounded assessment of the model’s performance.
Table 10: Overview of AI Benchmarks and Their Evaluation Focus
Benchmark Description
HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. [29]
MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. [30]
TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation. [31]
Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. [32]
GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. [33]
Base vs. Chat Models
Tab. 11 shows the comparative analysis of base and tuned models. Our findings suggest that: 1) Base models exhibit a stronger correlation with established benchmarks, implying that these models, in their fundamental form, possess a robust capacity for modeling language. This observation advocates for conducting initial evaluations using base models to capture their intrinsic language modeling capabilities. 2) The exploration of tuned models reveals potential deeper patterns in model tuning, suggesting that specific tuning strategies may unlock further enhancements in model performance across different tasks and domains.
Table 11: Correlation with Different Benchmarks, Base vs. Chat Models
Benchmarks Loss of Base Model Loss of Chat Model Delta
MMLU -0.61 -0.22 0.39
MMLU:Humanities -0.59 -0.42 0.17
MMLU:SocialSciences -0.54 0.42 0.96
MMLU:STEM -0.29 0.02 0.31
MMLU:Other -0.47 -0.1 0.37
Arc -0.46 0.34 0.8
Hellawasg -0.66 -0.57 0.09
Wnogrande -0.56 0.57 1.13
Longbench -0.04 -0.65 -0.61
GSM8K -0.4 0.22 0.62
TrustfulQA 0.05 0.39 0.34
Correlation alone Text Length
Tab. 12 is about a long text understanding benchmark Longbench.
For chat models in this task, the correlation between BPC and scores is logical—higher probabilities of generating the text correlate with higher scores, and this correlation strengthens with longer text lengths. This suggests a consistent alignment between the model’s language likelihood in long text and its scoring, reinforcing the chat model’s robustness in handling extensive text inputs.
However, an inverse trend is observed with base models, where a familiar text paradoxically results in lower scores on related comprehension tasks. This indicates that fine-tuning may alter the base models’ understanding of the texts, perhaps shifting their processing in ways that do not favor traditional comprehension metrics.
This indicates scenarios where a model scores high on language likelihood due to its text-generation skills but may not perform equally well on tasks that require deep semantic understanding or critical thinking. Thus, relying solely on language likelihood as an indicator of overall performance might overlook crucial aspects of cognitive and interpretative abilities that are essential for more complex applications.
Table 12: Correlation with Longbench values across different lengths
Length 10 30 50 100 200 300 400 500 800 1000 1200 1500 2000
Base Models 0.35 0.13 0.06 -0.03 -0.05 -0.06 -0.05 -0.05 -0.04 -0.04 -0.04 -0.04 -0.04
Chat Models 0.01 -0.07 -0.16 -0.39 -0.53 -0.58 -0.61 -0.63 -0.65 -0.65 -0.67 -0.67 -0.68
On the Genres of Corpora
As can be seen in Appendix  H, whether all models are observed or only those of a specific scale does not make a big difference. Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
Further analysis from Tab. 13 within the MMLU benchmark sections reveals interesting patterns. While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong. Specifically, STEM scores do not show strong correlation with STEM-related texts. Moreover, academic papers, even in Social Sciences and Humanities, exhibit stronger correlations than Internet data, which includes forums and news.
In this context, code-related data show similarities in performance correlation with academic texts, suggesting a closer alignment than previously anticipated. Similarly, Internet data, including forums and news, also shows comparative closeness in performance, indicating nuanced relationships across different types of content.
Table 13: Correlation Coefficients between BPC in specific areas and Domain-Specific Performance
Benchmark Academia STEM Academia non-STEM Code Internet Article Internet Chat
MMLU:STEM -0.17 -0.19 -0.11 -0.01 -0.01
MMLU:Humanities -0.54 -0.53 -0.51 -0.29 -0.39
MMLU:Social Sciences -0.49 -0.42 -0.46 -0.25 -0.31
Additionally, compared to chat forum texts, more formal academic texts demonstrate a stronger correlation with benchmarks. The genre of text has a minor impact on the sign of the correlation coefficient, which is relatively small compared to the differences across model types.
Appendix J Function Fitting Strategies
We considered several fitting strategies, each with its unique mathematical approach to model decay:
Linear Decay
Linear decay is represented by a straight line, suggesting a constant rate of decay over time. The model is given by the equation $y=ax+b$.
Exponential Decay
Exponential decay is modeled by an exponential function, expressed as $y=y_{0}\cdot e^{-kt}$, where $y_{0}$ is the initial value, $t$ is time or another decaying factor, and $k$ is the decay constant that denotes the rate of decay.
Power Law Decay
Power law decay implies a polynomial relationship and is expressed as $y=at^{-k}$, where $a$ is a positive constant, $t$ is time or another related factor, and $k$ represents the exponent characterizing the rate of decay.
Logarithmic Decay
Logarithmic decay suggests a decrease that becomes more gradual over time, typically modeled as $y=a+b\log(x)$, where $a$ and $b$ are constants, and $x$ represents time or another related factor.
The variance of the residuals for each model was calculated as follows:
Table 14: Variance of Residuals for Different Models
Model Variance
Linear 2.346e-3
Exponential 1.71e3
Power Law 3.78e-3
Logarithmic 3.33e-3
J.0.1 Findings
Our analysis found that the linear model provided the best fit, with the lowest variance among the residuals. This indicates a steady, unchanging rate of knowledge decay over time, as opposed to the more complex models that assume varying rates.
We aimed to evaluate the understanding and recency of knowledge across various language models, including GPT-4 and others. To accomplish this, we devised a methodology to calculate a weighted accuracy score for each model, reflecting not only its performance but also the contemporaneity of its knowledge base.
We began by assigning weights to different quarters, starting from 0 for the first quarter of 2015, up to 1 for the most recent quarter. This linear interpolation of weights was chosen to emphasize the value of more current information, underlining our hypothesis that accuracy concerning more recent knowledge is indicative of a model’s modernity and relevance.
The formula applied was straightforward: for each model, we multiplied its accuracy (Acc) by the corresponding weight based on the time difference since the first quarter of 2015. The resulting weighted accuracies were then summed up to yield an overall score for each model, which we used to rank them accordingly.
GPT-4 and other leading models demonstrated a notably later peak in knowledge, suggesting that their databases include more recent information. This is particularly valuable in fast-evolving fields where the timeliness of information is crucial.
Below is a partial presentation of rankings and scores for selected models, showcasing their weighted accuracy scores. The complete table can be found in Table 15.
Table 15: Model Scores
Model Name Score
GPT4-1106 794.275
GPT4-0613 699.8
Zephyr-7B-beta 601.35
InternLM-Chat-7B 592.15
GPT3.5-turbo-0613 578.025
Baichuan-13B-Chat 571.8
Qwen-14B-Chat 565.45
OPT-2.7B 547.375
Yi-6B-Chat 541.05
Mistral-7B-v0.1 528.725
Baichuan-7B-Chat 527.825
LLaMA-7B 524.075
Baichuan2-7B-Chat 517.725
Pythia-12B 513.375
Colossal-LLaMA2-7B-Base 513.35
LLaMA2-7B-Chat 508.4
LLaMA2-13B 508.15
Qwen-7B 506.7
Qwen-1.8B-Chat 505.225
Phi-2 497.4
Yi-6B 496.15
Zhongjing-Base 490.1
LLaMA2-7B 486.2
Falcon-rw-1B 481.15
ChatGLM3-6B 479.8
Baichuan2-7B-Base 479.175
Baichuan2-13B-Chat 463.75
OPT-13B 453.075
Baichuan2-13B-Base 446.225
Skywork-13B-Base 445.8
Qwen-1 8B 435.775
RWKV5-Eagle-7B 408.425
Qwen-7B-Chat 394.9
Phi-1.5 385.975
TinyLLaMA-1.1B-Chat-v0.6 366.125
Appendix K Details of Model Accuracy Peaks
Tab. 16 captures the peak performance periods of various models on Freshbench. It lists specific models along with their respective dates when they achieved the best performance.
Table 16: Detailed Record of Model Accuracy Peaks
Date Models
2015-01 Skywork-13B-Base, Phi-1.5
2015-07 Skywork-13B-Base, Phi-1.5
2016-01 Baichuan2-7B-Chat, Colossal-LLaMA-2-7B-Base, LLaMA-2-7B, Qwen-14B-Chat, Qwen-1.8B, Qwen-1.8B-Chat, Yi-6B, Yi-6B-Chat, Baichuan-13B-chat, Baichuan-7B-chat, ChatGLM3-6B, falcon-1B, InternLM-chat-7B, OPT-13B, Phi-2
2016-07 RWKV5-Eagle-7B
2017-01 GPT-3.5-turbo-0613
2017-07 Baichuan2-7B-Base
2018-01 Baichuan2-13B-Base
2018-07 Baichuan2-13B-Chat, LLaMA-2-13B-hf, Qwen-7B, GPT4-0613, GPT4-1106, LLaMA2-7B-Chat, LLaMA-7B, mistral-7B-v0.1, OPT-2.7B, Pythia-12B, Zephyr-7B-beta
2019-07 TinyLLaMA-1.1B-Chat
2020-01
2020-07
2021-01
2021-07
2022-01
2022-07
2023-01
2023-07
2024-01
2024-07
Appendix L Question Tag Details
Table 17: Detailed breakdown and sum of entries per category within each group of benchmarks.
Group Category Number of Entries
Political and Government Affairs Leader Entry/Exit 199
US Politics 338
Elections and Referenda 265
US Policy 223
Foreign Policy 242
Non-US Politics 415
Security and Conflict 413
Total: 2095
Economic and Business Items Business 575
Economic Indicators 253
Finance 284
Economic Policy 147
Total: 1259
Societal issues Items Society 379
Health 237
Environment 75
Sports 120
Entertainment 132
Total: 943
Science and Technology Items Technology 346
Total: 346
Appendix M Analysis of Best-Achieved Accuracy and Cutoff Time
Table 18: Analysis about the time of the best-achieved accuracy and its cutoff time.
Peak Period Model Peak Time of Accuracy Cutoff Time Interpolation
Post-2019 TinyLLaMA1.1B 2019-07 2023-11 52
2017 to 2018 GPT4-1106 2018-07 2023-06 59
LLaMA2-13B 2018-07 2023-02 55
Baichuan2-13B-Base 2018-01 2023-06 65
2015 to 2016 Baichuan2-7B-Chat 2016-01 2023-08 91
Colossal-LLaMA-2-7B-Base 2016-01 2023-09 92
LLaMA2-7B 2016-01 2023-02 85
Tab. 18 shows an analysis covering the period from 2015 to 2024, with the year 2019 serving as a pivotal midpoint, a distinct pattern emerges from the performance data of various language models: all models reached their peak accuracy before 2020. This observation points to a notable trend toward "Nostalgia", where models demonstrate optimal performance on data from the earlier part of the timeline analyzed. This tendency highlights the models’ alignment with the informational characteristics prevalent prior to the latter half of the observed period, suggesting their training data may disproportionately reflect earlier times.
The peak accuracy periods analyzed are based on the average accuracy of all questions within three-month intervals, to provide a stable and robust assessment of each model’s performance.
Appendix N Prediction Correlation over Time
The graphical representations in our analysis plot the changes in correlation coefficients over time. Each line in the graph represents the changing relationship between a category’s average adjusted performance and a specific benchmark, highlighting periods of stronger or weaker association.
Figure 7: Science and Technology Items - All Model Spearman Smoothed
Figure 8: Science and Technology Items - Base Model Spearman Smoothed
Figure 9: Societal Issues Items - All Model Spearman Smoothed
Figure 10: Political and Government Affairs - All Model Spearman Smoothed
Figure 11: Economic and Business Items - All Model Spearman Smoothed